OpenStudy (anonymous):

v=4/3pie r squared h solve for h (this is Geometry)

OpenStudy (anonymous):

\[\Large \begin{array}{l} V = \frac{4}{3}\pi {r^2}h\\ h = \frac{V}{{\frac{4}{3}\pi {r^2}}} = \frac{{3V}}{{4\pi {r^2}}} \end{array}\]

OpenStudy (anonymous):

Is that what you're asking?